# Decoding Generalization from Memorization in Deep Neural Networks

### v1

@@ 7 Discussion
In this work, we investigated the phenomenon of memorized networks not generalizing well, asking why the ability to generalize is apparently lost during the act of memorizing. We find, surprisingly, that the intrinsic ability to generalize remains present to a degree not previously recognized, and this ability can be decoded from the internals of the network by straightforward means.
@@
An interesting question is about why this phenomenon even occurs; naïvely one would expect that networks, on being trained with highly noisy data, discard the ability to generalize in favor of learning noise. Are there specific inductive biases that promote such generalization? And, do such mechanisms also promote generalization in networks whose training data isn’t corrupted significantly by such noise? It would also be instructive to study the dynamics of this form of generalization during training333which has been beyond the scope of this paper.. It is known (Arpit et al., 2017) that the model’s test accuracy transiently peaks in the early epochs of training with corrupted data, before dropping while training accuracy of the corrupted training data rises. It is unclear whether this transient rise in model generalization is caused by the subspace organization seen here, and if so, why such subspace organization isn’t degraded as much as the model’s test error over further epochs of training.
@@
The work has a number of implications. On the one-hand, it suggests that the ability to memorize and generalize may not be antithetical. Indeed, in multiple cases, we are able to construct single MASC classifiers that perform well both on the shuffled training labels as well as on the held-out test data that has true labels. Secondly, theories proposed to explain generalization in deep networks have traditionally argued for the setting where the data distribution is well-behaved, i.e. corresponding to real-world data, but not for data with shuffled labels. We suggest, in light of the present results, that such theories also ought to be able to explain why networks retain the ability to generalize even in the face of noisy training data. That is, a satisfactory understanding of generalization in deep networks should also cover the settings where the training data is noisy and its distribution is not well behaved. Thirdly and more pragmatically, techniques such as the MASC classifier might suggest a way of boosting generalization in trained Deep Networks, whose training data intrinsically contains varying degrees of label noise. While this has been beyond the scope of the present paper, possibilities of designing new techniques for learning subspaces that have good generalization ability could be explored. Indeed, it is possible that significantly better subspaces exist than the ones uncovered here, and it would be interesting to see how much the generalization accuracy can be improved by pursuing this direction. Relatedly, it is possible that other classifiers operating on layerwise outputs have better performance than MASC – a possibility that merits further exploration. Fourthly, it would be interesting to formulate a measure to study representational similarity between memorized & generalized networks to see if they use similar mechanisms. Does the answer depend on the particular class of networks (e.g. MLPs vs. CNNs)?
@@
Finally, the results here are reminiscent of a puzzling phenomenon observed in Neuroscience. In multiple settings (Shusterman et al., 2011; Miura et al., 2012; Stringer et al., 2021), in mice, rats and humans, it has been shown that a decoder using data from a subset of neurons from specific areas in the brain of a well-trained behaving animal has accuracy significantly better than the behavioral accuracy of the animal, even though the animal is motivated to do well on the task. It may therefore be that this is a phenomenon shared between brains and machines, whose underlying mechanisms and potential trade-offs remain to be investigated.
@@ Acknowledgments
Simran Ketha was supported by an APPCAIR Fellowship, from the Anuradha & Prashanth Palakurthi Centre for Artificial Intelligence Research. The work was supported in part by an Additional Competitive Research Grant from BITS to Venkatakrishnan Ramaswamy. The authors acknowledge the computing time provided on the High Performance Computing facility, Sharanga, at the Birla Institute of Technology and Science - Pilani, Hyderabad Campus.
Appendix A Appendix
A.1 Subspace creation
The process of creating subspaces, before using the MASC classifier, is shown schematically in Figure 4 for the MLP model.
Figure 4: Class-conditioned training data subspaces on layerwise outputs of MLP using PCA. Top: Schematic of MLP model used in the work. Bottom: Creating the class-conditioned training subspace for ReLU (128) layer where 128 are the number of neurons.
A.2 Model Details
The MLP model has 4 hidden layers with 128, 512, 2048 and 2048 units respectively. $ReLU$ activation was used after every layer and for classification $softmax$ activation was applied. We have trained the models with two different optimizers namely, $SGD$ and $Adam$. Learning rate of 1e-3 and momentum = 0.9 was used with $SGD$ optimizer. Learning rate = 1e-4 was used for $Adam$ in experiments. Batch size of 32 was used in all the models. Data set was normalized by dividing each pixel value with 255.
CNN network has 3 blocks, each consisting of two convolutional layers, one max pooling layer. These blocks are followed by three fully connected layers. Convolutional layers have 16, 32, and 64 filters, respectively with stride=1 and filter size = 3 × 3. Max pooling layer has stride of 1 and filter size of 2 × 2. The fully connected layers at the end has 250 units each. It was trained with $Adam$ optimizer with learning rate of 0.0002. For MNIST and Fashion-MNIST batch size of 32 whereas for CIFAR-10 batch size of 128 were used. Data set was normalized by subtracting the mean and diving by the standard deviation for each channel. $ReLU$ activation was used after every layer except pooling and $softmax$ activation for classification.
AlexNet model was slightly modified to for the use of each dataset. $Adam$ optimizer with learning rate of 0.0001 was used. For CIFAR-100, batch size of 128 and for Tiny ImageNet, batch size of 500 was used. All the results with respect to testing on AlexNet trained on Tiny ImageNet are shown with the validation dataset. CIFAR-100 dataset before training was normalized by subtracting the mean and diving by the standard deviation for each channel. No data normalization was performed on Tiny ImageNet dataset.
@@ A.2 Model Details
A.3 Training and testing performance of the models
Average training and testing accuracies of the models over three different runs used in this paper are shown in Tables 2 and 3.
Table 2: Average training accuracy in percentages of all the models over three runs over different corruption degrees (indicated in the last six columns). WD and WOD corresponds to with dropout and without dropout respectively.
@@ A.3 Training and testing performance of the models
CNN MNIST WOD 99.90 99.32 98.62 97.25 95.11 94.92
WD 99.90 95.73 87.28 77.70 71.18 57.24
Fashion-MNIST WOD 99.15 99.14 97.90 96.25 91.65 83.14
WD 99.13 94.89 85.33 76.45 67.55 58.81
CIFAR-10 WOD 99.70 99.29 99.26 99.03 99.02 39.69
WD 99.30 96.71 94.87 93.59 90.00 86.97
AlexNet CIFAR-100 WOD 99.19 99.15 99.11 99.16 99.14 97.88
WD 99.37 99.07 98.66 98.20 97.16 94.39
Tiny ImageNet WOD 99.92 99.90 99.91 99.93 87.71 85.95
Table 3: Average testing accuracy in percentages of all the models over three runs over different corruption degrees (indicated in the last six columns). WD and WOD corresponds to with dropout and without dropout respectively.
@@
CNN MNIST WOD 99.15 87.51 69.44 47.10 28.30 9.85
WD 99.13 86.19 55.99 37.37 25.85 9.95
Fashion-MNIST WOD 90.74 77.74 61.35 43.26 25.57 10.08
WD 91.36 74.56 53.20 34.92 24.32 9.68
CIFAR-10 WOD 74.95 60.48 46.15 30.96 18.32 9.89
WD 73.77 58.20 43.62 29.83 17.50 10.13
AlexNet CIFAR-100 WOD 36.75 28.44 20.53 9.64 3.43 0.96
WD 37.59 26.21 15.70 7.49 2.51 1.05
Tiny ImageNet WOD 15.88 9.74 5.44 2.02 0.73 0.43
A.4 Comparison of decoding results and early stopping test accuracies
Early stopping test model accuracy was added as a reference in this section to compare the results of testing accuracy of MASC on trained model with early stopping model accuracy. MASC accuracy over the layers of the network when the data is projected onto corrupted training subspaces is shown in Figure 5 and onto true training subspaces is shown in Figure 6. Best model testing accuracy and trained model testing accuracy are shown for reference. Best model testing accuracy corresponds to the accuracy of the testing data of the model if early stopping was used.
For MASC when the data is projected onto corrupted training subspace, in AlexNet-CIFAR-100, the MASC in at least one layer shows better performance than the best model testing accuracy for less than 60% corruption degree. For MLP-MNIST, the best model (early stopping) maintains over 90% accuracy even when the data is corrupted up to 80%. Despite the increase in corruptions (except 100% corruption), the accuracy of the last layer remains close to that of the best model accuracy. For CNN-FMNIST (except 100% corruption), in at least one layer MASC performance is near to that of best model testing accuracy.
@@ A.4 Comparison of decoding results and early stopping test accuracies
Figure 5: MASC test accuracy over the layers of the network when the data is projected onto corrupted training subspaces with the indicated corruption degree. Best Model Testing Accuracy corresponds accuracy of the testing data of the model if early stopping was used.
Figure 6: MASC test accuracy over the layers of the network when the data set is projected onto subspace corresponding to true training labels. Best Model Testing Accuracy corresponds accuracy of the testing data of the model if early stopping was used.

### v2

@@ 8 Discussion
In this work, we investigated the phenomenon of memorized networks not generalizing well, asking why the ability to generalize is apparently diminished due to the act of memorizing. We find, surprisingly, that the intrinsic ability to generalize remains present to a degree not previously recognized, and this ability can be decoded from the internals of the network by straightforward means. On the one hand, we design probes that use the subspace geometry of the corrupted training data to decode such better generalization to true labels. We also demonstrate using true labels post hoc that there exist subspaces that allow for an even more improved decoding. Furthermore, we show that such decoding can be leveraged to retrain the models to have better generalization to true labels. We also show (Appendix) that the internal representations of some deep networks trained on true labels, possess the ability to substantially memorize relabelings of its training data. Moreover, the new type of probe – MASC – that we use here, is relatively lightweight computationally while being easy to implement and lending itself to a straightforward geometric interpretation.
In building MASC, we were motivated by the manifold hypothesis in machine learning Goodfellow et al. (2016) that posits that high-dimensional data typically reside on a low-dimensional manifold. It has also been suggested (Brahma et al., 2015) that such manifolds in layerwise representations flatten across layers of deep networks. Fitting manifolds can be computationally expensive, so we were interested in examining the organization of classwise data in subspaces, even if such subspaces might be somewhat higher dimensional than the corresponding manifolds. Indeed, this view leads to the natural idea of classifying unseen data points by determining which class manifold it is closest to. MASC is simply a formalization of this idea. In particular, this classifier lends strong geometric motivation and intuition, in contrast to e.g. training a standard linear probe that iteratively minimizes a crossentropy loss. However, the reasons for the success of MASC in this setting are still largely unclear to us. The difficulty is that the principles that underlie the nature of layerwise representations in deep networks trained with standard techniques are not well understood at this time and it appears that such representations play a significant role in the success of MASC in the memorization setting. Indeed, it is even a bit surprising that the deep network does not directly leverage this structure to obtain better generalization to true labels, although that may also be because its loss function aims to maximize training accuracy which might run counter to the act of bettering generalization to true labels.
@@
An interesting question is about why this phenomenon even occurs; naïvely one would expect that deep networks, on being trained with noisy data, discard the ability to generalize in favor of learning noise. Are there specific inductive biases that promote such generalization to true labels? And, do such mechanisms also promote generalization to true labels in networks whose training data isn’t corrupted significantly by such noise? It would also be instructive to study the dynamics of this form of generalization to true labels during training. It is known (Arpit et al., 2017) that the model’s test accuracy transiently peaks in the early epochs of training with corrupted data, before dropping while training accuracy of the corrupted training data rises. It is unclear whether this transient rise in model generalization to true labels is caused by the subspace organization seen here, & if so, why such subspace organization isn’t degraded as much as the model’s test accuracy over further epochs of training. Additionally, certain deep networks trained on specific uncorrupted datasets seem to possess internal representations that are amenable to significant memorization, whereas others aren’t. The mechanistic basis of this ability is unclear & its possible connections to generalization to true labels in the such models merit further investigation.
It is interesting that probes other than MASC also often have generalization to true labels comparable to MASC, and this performance often manifests in layers different from those that have best accuracies of MASC. The reasons for this are unclear and remain to be investigated.
@@
The work has a number of implications. On the one-hand, it suggests that the ability to memorize and generalize may not be antithetical. Indeed, in multiple cases, we are able to construct single MASC classifiers that perform well both on the shuffled training labels as well as on the held-out test data that has true labels. Secondly, theories proposed to explain generalization to true labels in deep networks have traditionally argued for the setting where the data distribution is well-behaved, i.e. corresponding to real-world data, but not for data with shuffled labels. We suggest, in light of the present results, that such theories also ought to be able to explain why networks retain the ability to generalize even in the face of noisy training data. That is, a satisfactory understanding of generalization to true labels in deep networks should also cover settings where the training data is noisy and its distribution is not well-behaved. Thirdly and more pragmatically, techniques such as the MASC classifier might suggest a way of boosting generalization to true labels in trained deep networks, whose training data intrinsically contains varying degrees of label noise. While this has been beyond the scope of the present paper, possibilities of designing new techniques for learning subspaces that have good generalization to true labels could be explored. Indeed, it is possible that significantly better subspaces exist than the ones uncovered here, & it would be interesting to see how much the generalization to true labels accuracy can be improved by pursuing this direction. Relatedly, it is possible that other classifiers operating on layerwise outputs have better performance than MASC – a possibility that merits further exploration. Fourthly, it would be interesting to formulate a measure to study representational similarity between memorized & generalized networks to see if they use similar mechanisms. Does the answer depend on the particular class of networks (e.g. MLPs vs. CNNs)?
@@
Finally, the results here are reminiscent of a puzzling phenomenon observed in Neuroscience. In multiple settings (Miura et al., 2012; Stringer et al., 2021), in the rat olfactory system and the mouse visual system, it has been shown that a decoder using data from a subset of neurons from specific areas in the brain of a well-trained behaving animal has accuracy significantly better than the behavioral accuracy of the animal on novel trials, even though the animal is motivated to do well on the task. This implies that those animals have better innate generalization to true labels ability on that task – which can be easily decoded from a subset of their neurons – than is manifested by their behavior. It may therefore be that this is a phenomenon shared between brains and machines, whose underlying mechanisms and potential trade-offs remain to be investigated.
Broader impact statement
The Minimum Angle Subspace Classifier (MASC) is primarily an analytical probing method, and its direct societal risks are therefore limited. However, caution is advised when applying it to models trained on sensitive or proprietary data. In addition, probe-derived labels should not be repurposed for model retraining without appropriate safeguards, as this may reinforce existing biases or propagate systematic mislabeling; robust validation procedures, uncertainty thresholds, and human oversight are recommended. Finally, the diagnostic performance of MASC should not be interpreted as equivalent to the underlying model’s generalization to true labels ability. MASC’s results are about understanding representations and not about predicting how the model will perform when deployed.
@@ Acknowledgments
Simran Ketha was supported by an APPCAIR Fellowship, from the Anuradha & Prashanth Palakurthi Centre for Artificial Intelligence Research. The work was supported in part by an Additional Competitive Research Grant from BITS to Venkatakrishnan Ramaswamy. The authors acknowledge the computing time provided on the High Performance Computing facility, Sharanga, at the Birla Institute of Technology and Science - Pilani, Hyderabad Campus. We thank Harsha Varun Marisetty for compute assistance for training some of the models used here.
Appendix A Model details
Multilayer Perceptron (MLP) model has 4 hidden layers with 128, 512, 2048 and 2048 units respectively. $ReLU$ activation was used after every layer and for classification softmaxsoftmax activation was applied. We have trained the models with two different optimizers namely, $SGD$ and $Adam$. Learning rate of $1\text{\times}{10}^{-3}$ and momentum = 0.9 was used with $SGD$ optimizer. Learning rate = $1\text{\times}{10}^{-4}$ was used for $Adam$ in experiments. Batch size of 32 was used in all the models. The dataset was normalized by dividing each pixel value with 255.
Convolutional Neural Network (CNN) network has 3 blocks, each consisting of two convolutional layers, one max pooling layer. These blocks are followed by three fully connected layers. Convolutional layers have 16, 32, and 64 filters, respectively with stride=1 and filter size = 3 × 3. Max pooling layer has stride of 1 and filter size of 2 × 2. The fully connected layers at the end has 250 units each. It was trained with $Adam$ optimizer with learning rate of 0.0002. For MNIST and Fashion-MNIST batch size of 32 whereas for CIFAR-10 batch size of 128 were used. The dataset was normalized by subtracting the mean and diving by the standard deviation for each channel. $ReLU$ activation was used after every layer except pooling and softmaxsoftmax activation for classification.
AlexNet model was slightly modified for the use of each dataset. $Adam$ optimizer with learning rate of 0.0001 was used. For CIFAR-100, batch size of 128 and for Tiny ImageNet, batch size of 500 was used. All the results with respect to test on AlexNet trained on Tiny ImageNet are shown with the validation dataset. CIFAR-100 dataset before training was normalized by subtracting the mean and diving by the standard deviation for each channel. No data normalization was performed on Tiny ImageNet dataset.
ResNet-18 model was slightly modified for the use of CIFAR-10 dataset. $SGD$ optimizer with learning rate of 0.001 and momentum of 0.9 was used. Batch size of 32 was used for training. The dataset was normalized by subtracting the mean and diving by the standard deviation for each channel.
@@ Appendix A Model details
A summary of the models and datasets with training set size and number of parameters is in Table 2. The average training and test accuracies of the models over three different runs used in the paper are shown in Tables 3 and 4.
@@
ResNet-18 CIFAR-10 50,000 11,173,962
Table 3: Average training accuracy in percentages of all the models over three runs over different corruption degrees (indicated in the last six columns). The values are rounded to the second decimal place.
@@
CNN MNIST Adam 99.90 99.32 98.62 97.25 95.11 94.92
Fashion-MNIST Adam 99.15 99.14 97.90 96.25 91.65 83.14
CIFAR-10 Adam 99.70 99.29 99.26 99.03 99.02 39.69
AlexNet CIFAR-100 Adam 99.19 99.15 99.11 99.16 99.14 97.88
Tiny ImageNet Adam 99.92 99.90 99.91 99.93 87.71 85.95
ResNet-18 CIFAR-10 SGD 100.0 100.0 100.0 100.0 100.0 -
Table 4: Average test accuracy in percentages of all the models over three runs over different corruption degrees (indicated in the last six columns). The values are rounded to the second decimal place.
@@
CNN MNIST Adam 99.15 87.51 69.44 47.10 28.30 9.85
Fashion-MNIST Adam 90.74 77.74 61.35 43.26 25.57 10.08
CIFAR-10 Adam 74.95 60.48 46.15 30.96 18.32 9.89
AlexNet CIFAR-100 Adam 36.75 28.44 20.53 9.64 3.43 0.96
Tiny ImageNet Adam 15.88 9.74 5.44 2.02 0.73 0.43
ResNet-18 CIFAR-10 SGD 80.02 67.39 51.53 34.48 20.02 -
A.1 General terminology
The general terminology used in this work is as follows:
  • Model Training Accuracy: The accuracy of the model on the training set with corrupted labels.
  • Model Test Accuracy: The accuracy of the model on the test dataset with true labels.
  • Minimum Angle Subspace Classifier (MASC) Accuracy on Corrupted Training: Using Algorithm 1 (main paper), the accuracy of MASC predicted class labels with respect to corrupted labels of training dataset.
  • Minimum Angle Subspace Classifier (MASC) Accuracy on Original Training: Using Algorithm 1 (main paper), the accuracy of MASC predicted class labels with respect to true labels of training dataset.
  • Minimum Angle Subspace Classifier (MASC) Accuracy on Test: Using Algorithm 1 (main paper), the accuracy of MASC predicted class labels with respect to true labels of test dataset.
Appendix B Minimum Angle Subspace Classifier
A schematic illustrating the retraining process using Minimum Angle Subspace Classifier (MASC) is shown in Figure 9.
Figure 9: A schematic of the memorized model retraining pipeline that leverages MASC. Step 1: We first split the test dataset into 80%-20%. Step 2: We use the MASC accuracy on the corrupted subspaces on the 20% of the test dataset to identify the model’s best-layer. Step 3: Using the best-layer MASC predictions, we relabel the corrupted training dataset. Step 4: We train with the relabeled corrupted training dataset for upto 30 epochs and perform early stopping with patience of 3 by considering the 20% test dataset as a validation dataset. A similar process was followed while working with subspaces corresponding to true labels. Step 5: The test accuracy on the models is calculated with respect to the 80% test dataset.
Appendix C Generalization to true labels comparison: MASC vs. other probes
C.1 Experiment setup
For comparison, we evaluated five probes across all network layers: Logistic Regression (LR), K-Nearest Neighbors (KNN), Linear Discriminant Analysis (LDA), Quadratic Discriminant Analysis (QDA), and Nearest Class Mean (NCM).
For LR, we load the full corrupted training and test layer outputs along with labels, construct respectively data loaders with batch size 128, and train a PyTorch logistic regression model for 20 epochs using the $Adam$ optimizer (learning rate $1\text{\times}{10}^{-3}$) with cross-entropy loss. The test loader is used for inference.
For NCM, we load the corrupted training and test layer outputs along with labels, compute class-wise mean vectors from the training data, and classify each test example by cosine similarity to these class means.
For LDA and QDA, we load the same training and test layer outputs along with labels. We use scikit-learn library to build and train the corresponding models on the corrupted training layer outputs. Evaluating is performed using the test layer outputs.
For KNN, we use scikit-learn library to build the model with n_neighbors=5 and cosine distance, following the same training and evaluation protocol as LDA.
C.2 Results on additional models
Test accuracy of various probes over the layers of the MLP trained on CIFAR-10, CNN trained on MNIST and CIFAR-10, and AlexNet trained on Tiny ImageNet are shown in Figure 10. MASC test accuracy was overlaid for comparison.
Figure 10: Test accuracies for different probes over the layers of the network. Rows corresponds to plots which have the same corruption degree and the columns correspond to the models as noted. Test accuracy of the model and MASC are shown for comparison. FC corresponds to fully connected layer with ReLU activation whereas Flat corresponds to flatten layer without ReLU activation.
C.3 Computational cost across layers
The computational costs (in GFLOPs) for the various probes over the layers of the network are presented in Figure 11 for MLP–MNIST, CNN–Fashion-MNIST, and AlexNet–CIFAR-100, and in Figure 12 for the remaining models.
Figure 11: Computational costs (GFLOPS) for different probes over the layers of the network. Computational costs (GFLOPS) of MASC are shown for comparison. FC corresponds to fully connected layer with ReLU activation whereas Flat corresponds to flatten layer without ReLU activation.
Figure 12: Computational costs (GFLOPS) for different probes over the layers of the network. Computational costs (GFLOPS) of MASC are shown for comparison. FC corresponds to fully connected layer with ReLU activation whereas Flat corresponds to flatten layer without ReLU activation.
C.4 Comparing generalization to true labels of MASC with logistic regression and nearest class mean
Test accuracy of Logistic Regression (LR), Nearest Class Mean (NCM) and MASC over the layers of the network for MLP-MNIST, CNN-Fashion-MNIST, and AlexNet-CIFAR-100 are shown in Figure 13 and for MLP-CIFAR-10, CNN-MNIST, CNN-CIFAR-10, and AlexNet-Tiny ImageNet are shown in Figure 14.
The three probes exhibit distinct behaviors across network layers and corruption degrees. For AlexNet–CIFAR-100 and MLP–MNIST at 20% and 40% corruption, MASC outperforms both NCM and LR on most layers. In contrast, for CNN-based models, NCM generally achieves higher accuracy than MASC across the majority of layers, while MASC typically performs better than LR. However, there are instances where MASC matches or surpasses NCM – for example, in CNN–MNIST and CNN–Fashion-MNIST at 20%, 40%, and 60% corruption degrees, at the FLAT (576) layer.
Figure 13: Test accuracies for different probes over the layers of the network. Rows corresponds to plots which have the same corruption degree and the columns correspond to the models as noted. Test accuracy of the model and MASC are shown for comparison. FC corresponds to fully connected layer with ReLU activation whereas Flat corresponds to flatten layer without ReLU activation.
Figure 14: Test accuracies for different probes over the layers of the network. Rows corresponds to plots which have the same corruption degree and the columns correspond to the models as noted. Test accuracy of the model and MASC are shown for comparison. FC corresponds to fully connected layer with ReLU activation whereas Flat corresponds to flatten layer without ReLU activation.
Appendix D MASC comparison results with various PCA variance thresholds.
In this section, we compare MASC performance using subspaces constructed with different variance thresholds: 99%, 75%, 50%, 25%, and a single principal component (1PC). Results for MLP–MNIST, CNN–Fashion-MNIST, and AlexNet–CIFAR-100 are presented in Figure 15, with additional model results shown in Figure 16.
For these experiments, a single trained model was used, and the MASC computations for the variance comparison were repeated across different seeds. To assess robustness with respect to randomness, we performed five runs using seeds 10, 20, 30, 40, 50, and report the results as the mean along with the 95% confidence interval (CI).
For most models, we observed that under non-zero corruption degrees, several variance thresholds across many layers outperform the model’s generalization to true labels. However, no consistent pattern emerges across all model–dataset pairs as the variance thresholds vary.
For MLP–MNIST and CNN–Fashion-MNIST, subspaces capturing variance of 75%, 50%, 25%, and even 1-PC often outperformed the subspace capturing 99% variance across multiple layers. In contrast, for AlexNet–CIFAR-100, the 99% variance threshold performed better than the other thresholds at 20% and 40% corruption degrees. Additionally, we found that the 1-PC subspace exceeded the performance of other thresholds in several models.
@@ C.4 Comparing generalization to true labels of MASC with logistic regression and nearest class mean
Figure 15: Minimum Angle Subspace Classifier (MASC) test accuracy over the layers of the network when the data is projected onto corrupted training subspaces with the indicated corruption degree, for multiple models/datasets. Rows corresponds to plots with the same corruption degree & the columns correspond to the models, as noted. Test accuracy (dotted line) of the model is shown. FC corresponds to fully connected layer with ReLUReLU activation whereas Flat corresponds to flatten layer without ReLUReLU activation.
Figure 16: Minimum Angle Subspace Classifier (MASC) test accuracy over the layers of the network when the data is projected onto corrupted training subspaces with the indicated corruption degree, for multiple models/datasets. Rows corresponds to plots with the same corruption degree & the columns correspond to the models, as noted. Test accuracy (dotted line) of the model is shown. FC corresponds to fully connected layer with ReLUReLU activation whereas Flat corresponds to flatten layer without ReLUReLU activation.
Appendix E Computational costs of MASC
In this section, we report total time taken (in seconds) and computation cost (GFLOPS) of MASC for building the subspaces and inference over the layers of the network, for multiple models/datasets are shown in Figure 17 and Figure 18. The computational cost (GFLOPS) for building per-class subspaces are shown in Figure 19.
It was observed that, although inference [^11] By inference in the context of MASC, we mean act of using the already determined subspaces to classify incoming data point(s). requires more computational operations (in GFLOPs) than subspace construction, the total time taken for inference is nevertheless lower. This is likely because inference computations are highly optimized and benefit more from hardware acceleration than the covariance and PCA steps involved in building subspaces.
We also find a clear scaling trend with respect to layer dimensionality. Layers with higher feature dimensionality – such as FC(2048), FC(4096), or Flat(1024/576) – consistently require higher computational cost during subspace construction. In contrast, lower-dimensional layers (e.g., FC(128) or FC(250)) remain comparatively inexpensive. This behavior aligns with the expected quadratic dependence of PCA based subspace construction on the feature dimension, and highlights that computational cost grows steadily as the layer dimensionality increases.
Figure 17: Total time taken in seconds by MASC for building the subspaces and inference over the layers of the network, for multiple models/datasets. FC corresponds to fully connected layer with ReLUReLU activation whereas Flat corresponds to flatten layer without ReLUReLU activation.
Figure 18: Total computational cost (GFLOPS) by MASC for building the subspaces and inference over the layers of the network, for multiple models/datasets. FC corresponds to fully connected layer with ReLUReLU activation whereas Flat corresponds to flatten layer without ReLUReLU activation.
Figure 19: Computational cost (GFLOPS) by MASC for building per-class subspace over the layers of the network, for multiple models/datasets. FC corresponds to fully connected layer with ReLUReLU activation whereas Flat corresponds to flatten layer without ReLUReLU activation.